# A Bayesian statistical analysis of 897 pulsar flux density spectra

### v1

@@ III.2 Model selection with frequentist statistics
We also apply frequentist statistics, including the AIC with correction term, and robust regression using a Huber loss function to downweight outliers, following the approach of Jankowski_2018. The full procedure is described in Appendix A, where we also reproduce their dataset for direct comparison.
@@ IV.1 Comparison with previous works
Our analysis, based on a significantly expanded and thoroughly scrutinized dataset, leads to a substantive revision of the established understanding of pulsar radio spectra. The conventional wisdom, as suggested by Jankowski_2018, held that the simple power law was the dominant spectral shape, describing approximately 79% of pulsars. In stark contrast, we find that complex models exhibiting curvature or breaks are in fact the norm. This paradigm shift can be attributed to two key factors: the substantial improvement in data quality and quantity, and critical methodological refinements.
Jankowski_2018 analyzed 441 pulsars, with their conclusions partly constrained by the available data at the time. Our work leverages a dataset more than twice the size (897 pulsars), incorporating a wealth of new, high-quality measurements from major surveys such as Posselt_2023 and Spiewak_2022, which were not available previously. Crucially, our curation process excluded uncalibrated measurements, enhancing the overall reliability of the spectral fits.
A pivotal finding of our work is the identification of a methodological bias in the previous frequentist analysis. As demonstrated in Appendix A, the use of the corrected AIC by Jankowski_2018 inherently penalizes more complex models when the number of data points is small. For a pulsar with only 4 or 5 flux measurements, the AIC formula diverges for models with 3 or 4 parameters, effectively disqualifying them and artificially inflating the preference for the 2-parameter simple power law. When we reanalyze their dataset using the standard AIC, the prevalence of the simple power law in their own dataset drops from 64.1% to 26.8%, a figure that aligns far more closely with our new results. This indicates that the previously reported dominance of the simple power law was, to a significant extent, a statistical artifact.
@@
We also compared our findings with the more recent work of xu_statistical_2024, who reported that 68.51% of pulsars follow a simple power law. This discrepancy likely stems from their use of an outdated and less comprehensive version of the pulsar_spectra catalogue, which contained only 34 publications compared to the 156 in the version we employed. Our larger and more curated dataset, combined with our robust statistical treatment, provides a more complete and reliable picture of the pulsar population’s spectral characteristics.
@@ IV.2 Classification of millisecond pulsars
Following the definition of levin_high_2013, millisecond pulsars (MSPs) are those with period $P<70$ ms and spin-down rate $\dot{P}<10^{-17}~\mathrm{s}~\mathrm{s}^{-1}$. Periods and spin-down rates were obtained from the ATNF Pulsar Catalogue. Of the 897 pulsars analyzed in this work, 85 satisfy the MSP criteria. Their model classifications and log Bayes factor results are summarized in Tables 4 and 5.
The Bayesian model selection reveals that for 51.2% of MSPs (43 out of 84), the evidence decisively favors a curved or broken model over a simple power law ($\ln\mathrm{BF}\geq 5$). This finding directly contradicts the conclusion of Kuzmin_2001, who reported no evidence for low-frequency turn-overs in their study of 30 MSPs. The discrepancy likely arises from their smaller sample size and less sensitive statistical framework, which was unable to detect the subtler curvature that our larger dataset and Bayesian methodology can resolve.
@@
Best-fit model
$\ln\mathrm{BF}\geq 5$
vs. second-best model
$\ln\mathrm{BF}\geq 5$
vs. simple power law
@@ IV.3 Spectra peaked at gigahertz frequencies
GPS pulsars are characterized by a flux density maximum near 1 GHz. This behavior is commonly attributed to thermal free–free absorption associated with interaction between the pulsar and its surrounding environment (Kijak_2007; rajwade_gigahertz_2016; Kijak_2017). Our systematic search for GPS pulsars, which we define as sources whose flux density peaks between 0.6 and 2.0 GHz, has significantly expanded the known population. By applying stringent criteria that require both the observed data and the best-fit model to show a clear peak within this range, we have identified 74 confident GPS candidates.
A majority (58 out of 74) are new discoveries not listed in previous catalogues by kijak_low-frequency_2021 or xu_statistical_2024. An example is PSR J14136141 (Fig. 3), which shows a well-defined peak near 1 GHz in its data and broken power-law model. Conversely, we find that 37 previously proposed GPS pulsars do not meet our dual-requirement criteria. This discrepancy likely arises from irregular data sampling or outlier measurements that can bias model fits when not rigorously vetted.
@@
  • *: Pulsar previously identified as a GPS pulsar by kijak_low-frequency_2021.
  • †: Pulsar previously identified as a GPS pulsar by xu_statistical_2024.
Figure 4: Distribution of spectral indices for 121 pulsars best fit by the simple power law, and for 214 pulsars whose simple power law fits are not evidently worse than other models (any $\ln\mathrm{BF}\leq 3$).
Figure 5: Distribution of spectral indices for 20 MSPs and 101 non-MSPs best fit by the simple power law. MSPs exhibit a noticeably steeper mean spectral index.
Figure 6: Distributions of spectral indices for three best-fit categories: simple power law (121), low-frequency turn-over power law (110 pulsars), and high-frequency cut-off power law (47 pulsars).
@@ IV.4 Distribution of spectral indices
Our analysis confirms and refines the established distribution of pulsar spectral indices. For the 121 pulsars best fit by a simple power law, the spectral indices follow a Gaussian distribution with a mean of $1.61\pm 0.74$ (Fig. 4), consistent with previous studies (e.g. Jankowski_2018; bates_pulsar_2013). Furthermore, our analysis confirms the trend that MSPs have steeper spectral indices. For the 20 MSPs best fit by a simple power law, the mean spectral index is 1.94-1.94, notably steeper than the mean of 1.55-1.55 for non-MSPs in the same category (see Fig. 5).
Crucially, we find that the choice of spectral model significantly impacts the inferred spectral index. As shown in Fig. 6, the distributions for curved models are distinct. Pulsars fit with a low-frequency turn-over model have a steeper mean index of $2.45\pm 0.96$. Those fit with a high-frequency cut-off have a flatter mean index of $1.03\pm 1.01$.
@@ V Summary
Our principal conclusion is that the simple power law is not the dominant spectral form. Instead, complex spectra with curvature or breaks are the norm, with 68.8% of pulsars decisively better fit by such models. This paradigm shift, starkly contrasting with the previous consensus established by Jankowski_2018, is driven by our expanded dataset and the identification of a statistical bias in the previously used AIC metric, which artificially penalized complex models.
@@
    Millisecond pulsars are not spectrally simple; more than half exhibit decisive curvature, primarily described by broken power laws, and they possess steeper spectral indices than normal pulsars.
@@ Acknowledgements.
Appendix A Discussion on the AIC classification in Jankowski_2018
Jankowski_2018 published a dataset containing observations of 441 pulsars at three central frequencies: 728 MHz, 1382 MHz, and 3100 MHz. They also provided spectral classifications for these pulsars using an extended dataset that included additional measurements from the literature. In their analysis, the simple power law was reported as the best-fitting model for 62.6% of the 441 observed pulsars and for 79.1% of the 349 classified pulsars. The log-parabolic spectrum was the second most common model (35 pulsars). These results contributed to the widely adopted view that pulsar flux density spectra are generally well described by a simple power law, motivating extensive work on spectral index correlations based on this model.
To verify the classifications of Jankowski_2018, we reconstructed their dataset following the procedures described in their paper. Our reproduction draws from the ATNF Pulsar Catalogue (version 1.54, the same version as used in Ref. Jankowski_2018), flux density measurements reported in Jankowski_2018, and the additional literature sources they cited. A notable limitation is that, although Jankowski_2018 list multiple measurements at individual frequencies, the dataset included with their publication averages these into a single value per frequency. As a result, our reproduction necessarily reflects the available published data rather than the full multi-channel dataset. ATNF data were retrieved using the Python package psrqpy (pitkin_psrqpy_2018), and other literature values were obtained via the pulsar_spectra package (swainston_pulsar_spectra_2022). Missing entries, particularly from sieber_pulsar_1973 and maron_pulsar_2000, were manually added from the original publications.
Table 7: Classification of 441 pulsars according to table 5 of Jankowski_2018, compared with results obtained using our reproduced dataset. Excluding unclassified pulsars, the simple power law accounts for 79.1% of sources in Jankowski_2018 and 64.1% in our reproduction.
Model Classification in Jankowski_2018 Our reproduction
@@ Appendix A Discussion on the AIC classification in Jankowski_2018
Figure 8: Spectrum of PSR J1607-0032. Eight measurements with very small reported uncertainties from Posselt_2023 form a local turn-over feature that is inconsistent with the general trend.
We applied the frequentist goodness-of-fit method of Jankowski_2018 to our reproduced dataset. The fitting procedure used maximum-likelihood and least-squares methods implemented via the minuit C++ library and its Python interface iminuit (james_minuit_1975; dembinski_scikit_hepiminuit_2022), together with the Huber loss function and the “corrected AIC” statistic introduced by Jankowski_2018 and implemented in pulsar_spectra. We reproduced their fitting requirements as closely as possible. As shown in Table. 7, 202 of the 315 classified pulsars (64.1%) prefer the simple power law, in good agreement with Jankowski_2018. Although our dataset contains slightly fewer measurements because their multi-channel data were not available, the simple power law remains the most common best fit among the five models considered.
@@
To mitigate this structural disadvantage against (N1)(N-1)-parameter models when fitting $N$-point spectra, we recalculated the AIC without the correction term. Under this adjustment, the fraction of pulsars classified as simple power laws drops to 26.8%. In particular, 84 pulsars with only 4 or 5 measurements—previously forced into the simple power law category—shift to other models: 47 to the broken power law, 31 to the log-parabolic spectrum, 5 to the high-frequency cut-off power law, and 1 to the low-frequency turn-over model. This reclassification suggests that the dominance of the simple power law reported by Jankowski_2018 may be, at least in part, an artifact of their method penalizing multi-parameter models in sparsely sampled spectra.

### v2

@@ III.2 Model selection with frequentist statistics
We also apply frequentist statistics, including the AIC with correction term, and robust regression using a Huber loss function to downweight outliers, following the approach of Jankowski et al. (2018). The full procedure is described in Appendix A, where we also reproduce their dataset for direct comparison.
@@ IV.1 Comparison with previous works
Our analysis, based on a significantly expanded and thoroughly scrutinized dataset, leads to a substantive revision of the established understanding of pulsar radio spectra. The conventional wisdom, as suggested by Jankowski et al. (2018), held that the simple power law was the dominant spectral shape, describing approximately 79% of pulsars. In stark contrast, we find that complex models exhibiting curvature or breaks are in fact the norm. This paradigm shift can be attributed to two key factors: the substantial improvement in data quality and quantity, and critical methodological refinements.
Jankowski et al. (2018) analyzed 441 pulsars, with their conclusions partly constrained by the available data at the time. Our work leverages a dataset more than twice the size (897 pulsars), incorporating a wealth of new, high-quality measurements from major surveys such as Posselt et al. (2023) and Spiewak et al. (2022), which were not available previously. Crucially, our curation process excluded uncalibrated measurements, enhancing the overall reliability of the spectral fits.
A pivotal finding of our work is the identification of a methodological bias in the previous frequentist analysis. As demonstrated in Appendix A, the use of the corrected AIC by Jankowski et al. (2018) inherently penalizes more complex models when the number of data points is small. For a pulsar with only 4 or 5 flux measurements, the AIC formula diverges for models with 3 or 4 parameters, effectively disqualifying them and artificially inflating the preference for the 2-parameter simple power law. When we reanalyze their dataset using the standard AIC, the prevalence of the simple power law in their own dataset drops from 64.1% to 26.8%, a figure that aligns far more closely with our new results. This indicates that the previously reported dominance of the simple power law was, to a significant extent, a statistical artifact.
@@
We also compared our findings with the more recent work of Xu et al. (2024), who reported that 68.51% of pulsars follow a simple power law. This discrepancy likely stems from their use of an outdated and less comprehensive version of the pulsar_spectra catalogue, which contained only 34 publications compared to the 156 in the version we employed. Our larger and more curated dataset, combined with our robust statistical treatment, provides a more complete and reliable picture of the pulsar population’s spectral characteristics.
@@ IV.2 Classification of millisecond pulsars
Following the definition of Levin et al. (2013), millisecond pulsars (MSPs) are those with period $P<70$ ms and spin-down rate $\dot{P}<10^{-17}~\mathrm{s}~\mathrm{s}^{-1}$. Periods and spin-down rates were obtained from the ATNF Pulsar Catalogue. Of the 897 pulsars analyzed in this work, 85 satisfy the MSP criteria. Their model classifications and log Bayes factor results are summarized in Tables 4 and 5.
The Bayesian model selection reveals that for 51.2% of MSPs (43 out of 84), the evidence decisively favors a curved or broken model over a simple power law ($\ln\mathrm{BF}\geq 5$). This finding directly contradicts the conclusion of Kuzmin and Losovsky (2001), who reported no evidence for low-frequency turn-overs in their study of 30 MSPs. The discrepancy likely arises from their smaller sample size and less sensitive statistical framework, which was unable to detect the subtler curvature that our larger dataset and Bayesian methodology can resolve.
@@
Best-fit model $\ln\mathrm{BF}\geq 5$ vs. second-best model $\ln\mathrm{BF}\geq 5$ vs. simple power law
@@ IV.3 Spectra peaked at gigahertz frequencies
GPS pulsars are characterized by a flux density maximum near 1 GHz. This behavior is commonly attributed to thermal free–free absorption associated with interaction between the pulsar and its surrounding environment (Kijak et al., 2007; Rajwade et al., 2016; Kijak et al., 2017). Our systematic search for GPS pulsars, which we define as sources whose flux density peaks between 0.6 and 2.0 GHz, has significantly expanded the known population. By applying stringent criteria that require both the observed data and the best-fit model to show a clear peak within this range, we have identified 74 confident GPS candidates.
A majority (58 out of 74) are new discoveries not listed in previous catalogues by Kijak et al. (2021) or Xu et al. (2024). An example is PSR J14136141 (Fig. 3), which shows a well-defined peak near 1 GHz in its data and broken power-law model. Conversely, we find that 37 previously proposed GPS pulsars do not meet our dual-requirement criteria. This discrepancy likely arises from irregular data sampling or outlier measurements that can bias model fits when not rigorously vetted.
@@
  • *: Pulsar previously identified as a GPS pulsar by Kijak et al. (2021).
  • †: Pulsar previously identified as a GPS pulsar by Xu et al. (2024).
Figure 4: Distribution of mean spectral indices for 121 pulsars best fit by the simple power law, and for 214 pulsars whose simple power law fits are not evidently worse than other models (any $\ln\mathrm{BF}\leq 3$).
Figure 5: Distribution of mean spectral indices for 20 MSPs and 101 non-MSPs best fit by the simple power law. MSPs are steeper, but the difference is small ($\lesssim 1\sigma$).
Figure 6: Distributions of mean spectral indices for three best-fit categories: simple power law (121), low-frequency turn-over power law (110 pulsars), and high-frequency cut-off power law (47 pulsars).
@@ IV.4 Distribution of spectral indices
Our analysis confirms and refines the established distribution of pulsar spectral indices. For the 121 pulsars best fit by a simple power law, the spectral indices follow a Gaussian distribution with a mean of 1.61-1.61 and a standard deviation of $0.74$ (Fig. 4), consistent with previous studies (e.g. Jankowski et al., 2018; Bates et al., 2013). Furthermore, our analysis confirms a modest trend toward steeper spectral indices for MSPs. For the 20 MSPs best fit by a simple power law, the mean spectral index is 1.94-1.94, compared with 1.55-1.55 for non-MSPs in the same category (see Fig. 5).
Crucially, we find that the choice of spectral model significantly impacts the inferred spectral index. As shown in Fig. 6, the distributions for curved models are distinct. Pulsars fit with a low-frequency turn-over model have a steeper mean index of 2.45-2.45 and a standard deviation of $0.96$. Those fit with a high-frequency cut-off have a flatter mean index of 1.03-1.03 and a standard deviation of $1.01$.
@@ V Summary
Our principal conclusion is that the simple power law is not the dominant spectral form. Instead, complex spectra with curvature or breaks are the norm, with 68.8% of pulsars decisively better fit by such models. This paradigm shift, starkly contrasting with the previous consensus established by Jankowski et al. (2018), is driven by our expanded dataset and the identification of a statistical bias in the previously used AIC metric, which artificially penalized complex models.
@@
    Millisecond pulsars are not spectrally simple; more than half exhibit decisive curvature, primarily described by broken power laws, and they show a slightly steeper mean spectral index than normal pulsars.
@@ Acknowledgements.
Appendix A Discussion on the AIC classification in Jankowski et al. (2018)
Jankowski et al. (2018) published a dataset containing observations of 441 pulsars at three central frequencies: 728 MHz, 1382 MHz, and 3100 MHz. They also provided spectral classifications for these pulsars using an extended dataset that included additional measurements from the literature. In their analysis, the simple power law was reported as the best-fitting model for 62.6% of the 441 observed pulsars and for 79.1% of the 349 classified pulsars. The log-parabolic spectrum was the second most common model (35 pulsars). These results contributed to the widely adopted view that pulsar flux density spectra are generally well described by a simple power law, motivating extensive work on spectral index correlations based on this model.
To verify the classifications of Jankowski et al. (2018), we reconstructed their dataset following the procedures described in their paper. Our reproduction draws from the ATNF Pulsar Catalogue (version 1.54, the same version as used in Ref. Jankowski et al. (2018)), flux density measurements reported in Jankowski et al. (2018), and the additional literature sources they cited. A notable limitation is that, although Jankowski et al. (2018) list multiple measurements at individual frequencies, the dataset included with their publication averages these into a single value per frequency. As a result, our reproduction necessarily reflects the available published data rather than the full multi-channel dataset. ATNF data were retrieved using the Python package psrqpy (Pitkin, 2018), and other literature values were obtained via the pulsar_spectra package (Swainston et al., 2022). Missing entries, particularly from Sieber (1973) and Maron et al. (2000), were manually added from the original publications.
Table 7: Classification of 441 pulsars according to table 5 of Jankowski et al. (2018), compared with results obtained using our reproduced dataset. Excluding unclassified pulsars, the simple power law accounts for 79.1% of sources in Jankowski et al. (2018) and 64.1% in our reproduction.
Model Classification in Jankowski et al. (2018) Our reproduction
@@
Figure 8: Spectrum of PSR J1607-0032. Eight measurements with very small reported uncertainties from Posselt et al. (2023) form a local turn-over feature that is inconsistent with the general trend.
We applied the frequentist goodness-of-fit method of Jankowski et al. (2018) to our reproduced dataset. The fitting procedure used maximum-likelihood and least-squares methods implemented via the minuit C++ library and its Python interface iminuit (James and Roos, 1975; Dembinski et al., 2022), together with the Huber loss function and the “corrected AIC” statistic introduced by Jankowski et al. (2018) and implemented in pulsar_spectra. We reproduced their fitting requirements as closely as possible. As shown in Table. 7, 202 of the 315 classified pulsars (64.1%) prefer the simple power law, in good agreement with Jankowski et al. (2018). Although our dataset contains slightly fewer measurements because their multi-channel data were not available, the simple power law remains the most common best fit among the five models considered.
@@
To mitigate this structural disadvantage against (N1)(N-1)-parameter models when fitting $N$-point spectra, we recalculated the AIC without the correction term. Under this adjustment, the fraction of pulsars classified as simple power laws drops to 26.8%. In particular, 84 pulsars with only 4 or 5 measurements—previously forced into the simple power law category—shift to other models: 47 to the broken power law, 31 to the log-parabolic spectrum, 5 to the high-frequency cut-off power law, and 1 to the low-frequency turn-over model. This reclassification suggests that the dominance of the simple power law reported by Jankowski et al. (2018) may be, at least in part, an artifact of their method penalizing multi-parameter models in sparsely sampled spectra.